OpenStudy (anonymous):

How many moles of ethane (C2H6) would be needed to react with 62.3 grams of oxygen gas?

OpenStudy (anonymous):

@Hayhayz

Hayhayz (hayhayz):

oh boy.. this is what I'm doing too and its confusing for me.. are you by any chance in flvs?

OpenStudy (anonymous):

Nope.

OpenStudy (photon336):

You need to write out the balanced equation first. \[2C_{2}H_{6}+7O_{2} \rightarrow 4CO_{2}+6H_{2}O\]

OpenStudy (photon336):

what this tells us is that 2 moles of ethane react with 7 moles of oxygen gas. we are given 62.3 grams of Oxygen so we need to convert this to moles. We do this by \[(\frac{ mol }{ 32grams_{oxygen gas} })*(62.3grams) =1.9468 =~ 2 moles_{oxygen gas}\]

OpenStudy (photon336):

now that we have 2 moles of oxygen gas we need to figure out how many moles of ethane we need so we multiply this number by the molar ratio. do you know how to do this?

OpenStudy (anonymous):

no..

OpenStudy (photon336):

We need to multiply the number of moles of oxygen gas by the molar ratio of oxygen to ethane. \[\frac{ 2C_{2}H_{6} }{ 7O_{2} }*2mol_{Oxygen gas} = 0.57 moles, ethane\]

OpenStudy (anonymous):

That's it..?

OpenStudy (photon336):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (photon336):

no problem welcome to Open Study